LHC 2010 - latest news

Week 45: Machine coordinators: Mike Lamont –Ralph Assmann

Main aims of the week:

Outline of ion commissioning plan

LHC coordination vistar



Link to report from morning meeting

Monday 15th November

  • 00h30: Beam dump.
  • 01h30: Injection. 121b x 121b.
  • 07h42: Stable beams. Fill #1492. Reach 2.5e25 cm-2 s-1.

Slides from 8:30 meeting

Sunday 14th November

  • 00h32: Stable beams. Fill #1490. 121b x 121b. Exceed 2e25 cm-2 s-1 luminosity.
  • 06h26: Lost QPS-OK in Q9.L7 (SEU suspected). Same location as yesterday but different problem: FIP communication. This type of problem occurred yesterday in Q9.R7 and is not addressed by firmware upgrade deployed during Saturday day accesses.
  • 08h05: Beam dump due to power interlock in L5. Ramp down.
  • 09h00: Accesses for QPS problem in L7 and switch of RQTL11_L5B1 board.
  • 12h05: Recovery from access.
  • 13h53: Injection. Settings for reduced transverse blow-up in the PS, that should give smaller emittance (3 -> 2.1 – 2.8 um).
  • 16h49: Start ramp. 121 x 121 bunches. Long. blow up used to keep bunch length at or above 1ns.
  • 18h05: Stable beams. Fill #1491. 121b x 121b. 

Slides from 9:00 meeting

Saturday 13th November

  • 01h04: Stable beams. Fill #1489. 69b x 69b. On IR7L and IR7R, each one QPS equipment tripped (SEU).
  • 10h44: Beam dump by request. Preparing QPS accesses to reset QPS equipment L7 and R7. Decide to go for firmware upgrade. Will take some hours.
  • 11h45: Machine prepared for accesses: Sectors switched off and locked: 45, 56, 67, 78, 81. Vac valves closed in UJ67 and UJ83.
  • 11h52: Access started in point 6
  • 12h16: In response to request from the tunnel, TI asked to switch on lights in the half octant left of IR7. ODH and evacuation alarms triggered by lights. Tunnel sections evacuated. Fire brigade into tunnel.
  • 19h25: Firmware upgrade to improve the tolerance to SEUs completed. Closing the LHC... we leave the lights on.
  • 21h40: Injection 121 on 121. Long – around an hour. BPM for dump checked and OK.

Slides from 9:00 meeting

Friday 12th November

  • 00h53: Stable beams. 68 on 69.
  • 06h41: trip of RQD.A23
  • 06h49: trip of sector 67 when loading the rampdown start BP (SEU?).
  • 09h49: Access for QPS problems. See Saturday report for details. One problem likely related to SEU (single event upset).
  • 19h04: QPS accesses finished. Starting recovery.
  • 20h52: Injection. RF voltage modulation setup.
  • 22h42: Start ramp. 69b x 69b.

Slides from 8:30 meeting

Thursday 11th November

  • 00h03: Recover and precycle.
  • 00h56: Trip RQT13.L2B2. Piquet. Precycle.
  • 03h14: Injection. Missing frequently injection because of wrong bucket. During the injection for physics, we got one train shifted by one RF bucket not intercepted by the SPS BQM. So we dumped and refill.
  • 05h24: Start ramp. BCT logging stopped working. 69b. Low intensity.
  • 05h32: Beam dump. RB_ENERGY_CHECK SIS interlock. Same event as in the morning of the 27th of October. XPOC failed also because the energy reading is -1 GeV.
  • 05h56: Trip again RCBXH1.R2. Access.
  • 07h08: RCBXH1.R2 fixed. Board bypassed. Work through access items from standby list. Access planned until 11h00.
  • Problems during access: lights on à ODH alarm, patrol lost, problem in 81, PP60A (earth fault in a 18 Kv cell à PC problem), lights off à ODH alarm, various trips, …
  • 16h35: Access finished. Recover and precycle.
  • 18h17: Ready for injection.
  • 18h23: No beam from injectors: LEIR/PS problem. The beam is not extracted. LHC had mastership.
  • 20h08: Injection.
  • 21h00: Beam presence flag problem.
  • 21h30: Timing missing
  • 23h50: Start ramp

Slides from 8:30 meeting

Wednesday 10th November

  • 12h30: Adjust mode for end-of-fill study.
  • 13h21: Rampdown and precycle.
  • 14h13: Problem with timing card in RF front end at SPS.
  • 14h39: PS RF cavity down now.
  • 15h10: Injection.
  • 15h45: Tested dynamic change of reference for the Tune-FB and fitter window.
  • 16h19: PS RF problem.
  • 17h26: Injection.
  • 18h28: Start ramp. 69 bunches.
  • 18h41: RCBXH1.R2 tripped, beam dump.
  • 21h31: Recover and precycle.
  • 21h48: RCBXH1.R2 tripped again. Access. Connector? Old card.

Slides from 8:30 meeting

Tuesday 9th November

  • 01h01: Stable beams. Fill #1483.
  • 10h00: Beam dump. Ramp down and precycle.
  • Injection studies with 4 x 500 ns from SPS
  • Test ramp - new collision beam process
  • 22h50: Stable beams with 17 bunches, 16 colliding pairs fill# 1485 - initial lumi 3.5e24 cm-2s-1.

Slides from 8h30 meeting

Monday 8th November

  • 8h43: Injection.
  • 9h52: Start ramp. 2b of 1e10 per beam.
  • 11h20: Stable beams. First for ions. Fill #1482.
  • 12h57: Retracted TCTVB’s in IR2 to +- 23 mm (manually). Avoids shadowing problem for ZDC.
  • 14h19: Some orbit optimization: explains collimation issues observed?
  • 19h31: Adjust handshake starting.
  • 20h00 in adjust. End of fill loss map for reproducibility check. Loss map with TCTVB’s in IR2 back in.
  • 20h30: Rampdown and precycle. RF adjustments.
  • 22h20: Injection. Trip of RF module 1B1. RF adjustments.
  • 23h47: Start ramp. 5 bunches of 1e10 per beam.

Slides from 8:30 meeting